OpenStudy (anonymous):

need help with stating excluded values for a rational expression. m^2-1 over 2m^2-m-1

OpenStudy (anonymous):

I have a similar problem, this stuff seems fishy, but I know it must be simple

OpenStudy (anonymous):

can you figure this out?

OpenStudy (amistre64):

if we factor the top and bottom, we can determine the "excluded" values...

OpenStudy (amistre64):

m^2-1 factors to: (m-1)(m+1) its called the difference of squares 2m^2-m-1 factors to: (2m-3)(m+1) so we have: (m-1)(m+1) ----------- anything that makes the bottom a "zero" is excluded (2m-3)(m+1) m=3/2 and m=-1 should be illegal to use.. at m=3/2 we have a vertical asymptote; and at m=-1 we have a "hole" in the graph.